# Through the Lens of Human-Human Collaboration: A Configurable Research Platform for Exploring Human-Agent Collaboration

### v1

@@ 2.3. Human-Human Remote Collaboration as an Analytical Lens of Exploring Human-Agent Collaboration
LLM agents can be viewed as partially analogous to remote human collaborators: both rely on structured and explicit information channels, such as text or voice, and lack access to non-verbal cues (e.g., facial expressions, tone) that shape face-to-face interaction (Short et al., 1976; Walther, 1996; Olson and Olson, 2000). Early CSCW research on remote collaboration has identified key challenges and interaction variables that shape distributed teamwork. With respect to information channels, for instance, Media Richness Theory explains why low-bandwidth channels carry less contextual information and make negotiation harder (Daft and Lengel, 1986), with several experiments demonstrating that “leaner” media like text are less effective than “richer” media like video for conveying ambiguity and building consensus (Short et al., 1976; Straus, 1997). These results are critical to humans-agent collaboration because the interactions primarily happened via text. Interface cues such as typing indicators, meta messages, could enrich these channels and influence how users interpret agents’ intent and actions (Erickson and Kellogg, 2000; Iftikhar et al., 2023).
@@
Another recurring theme is the “mutual knowledge problem”, which arises when collaborators lack “common ground” (Clark and Brennan, 1991). Without shared references and assumptions, people struggle to know what their partners know, leading to misunderstandings and flawed assumptions (Cramton, 2001). Interfaces that add timely acknowledgments and visible states of collaborators could reduce these failures (Kraut et al., 2003). Research on workspace awareness further strengthened this idea by defining cues about presence, activity, and intent that help people predict collaborators’ actions (Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002). For LLM agents, this can be achieved by asking agents to generate future plans with rationales that could be made visible to human collaborators through interface designs. Task structure and coupling also matter. Olson and Olson (2000) showed why “distance matters”, highlighting how timing, handoffs, and availability complicate remote collaboration. Hinds and Bailey (2003) also analyzes how distributed teams coordinate under variable latency and partial information.
@@
In particular, a variety of lab-based experiments were designed to systematically analyze the impact of different interaction variables in a controlled setting. For instance, in DayTrader, researchers manipulate the communication medium during negotiation in a social-dilemma setting and identified that the choice of medium shifts trust formation and deal quality (Bos et al., 2002). Bos et al. (2004) further illustrated trading biases and subgroup formation caused by the differences between in-group (collocated partners who can talk face-to-face) and out-group (remote collaborators) in the Shape Factory experiment. NeoCITIES imposes time pressure and role interdependence in a shared information space (McNeese et al., 2005; Schelble et al., 2022), whereas Passcode examines how teams form and update shared mental models during problem solving when only partial clues are available (Gero et al., 2020). Despite LLM agents may differ from human remote collaborators in how they explain their decisions and how stable their identity recognition and memory are, research in the explainability of AI has pointed to feasible approaches by enabling provenance, disclosure of uncertainties, and human-understandable rationales given LLMs’ strong reasoning and text generation capabilities (Liao et al., 2020; Hoffman et al., 2023).
Our research platform draws on this profound line of CSCW work by enabling the manipulation of different interaction variables on the interface designs through the re-implementation of conventional experimental paradigms, but in the context of human-agent collaboration.
@@ 2.4. Research-Focused Toolkits and Platforms for Collaboration Studies
A growing number of research toolkits or platforms were designed to facilitate collaboration studies (Fig 2). We group them into four families: groupware/toolkits, general-purpose platforms, game-based research environments, along with newer human–AI collaboration systems and agent frameworks.
Early groupware toolkits in the HCI community provided fundamental building blocks for synchronous, distributed collaborative tasks between humans, which focused on enabling shared views, real-time session management, and widgets for customization, as seen in Share (Greenberg, 1990) and GroupKit (Roseman and Greenberg, 1996; Greenberg and Roseman, 1999). Researchers reflected on the evaluation of such toolkits with strategies outlined for evidence, generality, and adoption (Ledo et al., 2018). There exist other general-purpose online lab platforms that aim to lower the barrier for synchronous studies at scale. For instance, LabintheWild (Reinecke and Gajos, 2015) supports public-facing HCI experiments with built-in participant recruitment support and data analytics. Empirica (Almaatouq et al., 2021) provides real-time grouping, messaging, and logging for interactive tasks; oTree (Chen et al., 2016) and nodeGame (Balietti, 2017) implement economic games and behavioral experiments with session management and data capture. Nevertheless, these platforms only focus on human studies with no support for agent integration, and do not support manipulation of collaboration variables through interface customizations. In the HCI community, a number of studies also explored the usability of game environments as research platforms. Such game-based environments emphasize tightly-coupled coordination and performance, such as the real-time collaborative cooking game Overcooked (Carroll et al., 2019); researchers also leveraged Left 4 Dead 2 to compare perceptions of human versus NPC teammates during cooperative play (Wehbe et al., 2017).
@@
On the other hand, a large number of research platforms were proposed for developing and evaluating LLM agents. Multi-agent frameworks compose LLM agents and tools for task automation or agent–agent interaction, such as AutoGen (Wu et al., 2024), AgentVerse (Chen et al., 2023), and CAMEL (Li et al., 2023). Another family of agent-based frameworks, such as WebArena (Zhou et al., 2023) and Melting Pot (Leibo et al., 2021), was designed to systematically assess agent performances and social behaviors in a simulated environment. Moreover, agent-based social simulacra demonstrate LLM agents’ capacity to model social and cognitive behaviors at scale in a sandbox environment (Park et al., 2023, 2024) or support large-scale simulation with thousands of LLM agents interacting in a spatial or societal environment (Piao et al., 2025). Nevertheless, these agent-oriented frameworks and simulation platforms emphasize agent performance, simulation fidelity, or agent–agent dynamics, not for human-agent interactions.
Our work fills the critical gap between the rich theoretical heritage of CSCW and the practical needs of modern human-agent interaction research. Existing platforms either emphasize human collaboration studies with no support for agent integration or, on the other hand, focus on multi-agent interactions and simulations with no human engagement; besides, there is a lack of research platforms that provide researchers the rigorous controls needed for human-agent collaboration research. One recently proposed field-experiment platform called Pairit shares a similar conceptual idea with our platform, in which researchers compare humans’ performance in a human-human setting and a human-agent setting for a co-edit task (Ju and Aral, 2025), but this platform lacks customizability for different experimental paradigms and interface designs. Unlike existing platforms, our proposed platform is explicitly designed for the manipulation of theory-driven collaboration variables, and it is extensible and configurable for HCI researchers to re-implement, adapt, and invent new paradigms beyond conventional experimental paradigms used in human collaboration research. More importantly, our platform is purpose-built for human-agent collaboration, with a standardized agent integration protocol for agent integration and a data logging module for collecting both human behavior and agent reasoning.
@@ 3. System Design
To address the need for reproducible and extensible research on human-agent collaboration, we designed and implemented an open-source, fully configurable research platform. It serves as a flexible infrastructure for controlled experiments by balancing ease-of-use with customizability. The platform’s architecture encompasses four components: Participant Interface for user interaction, Researcher Interface for experiment management, Experiment Controller that serves as the backend engine, and an Agent Context Protocol (ACP) to standardize agent integration.
Researchers can adapt our platform for different experimental paradigms in two ways: through a declarative Experiment Configuration Language (ECL) (example provided in Appendix B) that defines all entities, actions, and interface modules, or by directly extending the open-source codebase. This dual approach supports seamless adaptation of existing paradigms, such as social dilemma games (Bos et al., 2002) and collaborative decision-making tasks (Zheng et al., 2023b), and the development of entirely new ones. Researchers can adjust the experiment parameters and manipulate the interaction controls for different conditions through the researcher interface. Our design is guided by four principles derived from foundational CSCW and HCI research: generality and extensibility, modularity, theory-grounded customizability, and standardized agent integration.
We address open-science needs by releasing source code (upon acceptance) and by supporting the creation, validation, and sharing of configuration templates, expecting our platform to support the broader research community in human-agent collaboration research.
@@ 3.1. Design Principles and Rationale
The platform targets synchronous or near-synchronous, interface-mediated collaboration among multiple participants that may be all human, all LLM agents, or mixed human-agent teams. Four principles guide the design of our research platform, each of which is grounded in foundational HCI work.
The first principle is generality and extensibility across experimental paradigms. Instead of hard-coding the platform’s functionalities for a single experiment, we abstract task, resource, and participant models to a configuration framework that separates classes from instances. The configuration language was thus designed to adhere to this framework. With the configuration language, researchers can either re-use and customize existing classes or instantiate new classes for experiment-specific artifacts and visual elements, and define their custom attributes correspondingly. This object-centered design supports resource negotiation and exchange, information sharing, and collaborative decision making within a unified architecture. As specified in Bos et al. (2004), most experiment paradigms are “not intended to be a full simulation of real-world collaboration,” but focus on reproducing certain collaborative dynamics.
The second principle is the modularization of interfaces and logic. Modularization has been repeatedly identified by prior research as essential for adaptable and replicable experimental platforms (Greenberg and Roseman, 1999; Gutwin and Greenberg, 2002). In our platform, the participant and researcher interfaces are composed of independent modules whose visibility and parameters. The participant interface consists of separate modules for private states, social interactions, task operations, and shared awareness that can be independently toggled and customized. The researcher interface consists of four primary modules: experiment selection, session configuration, real-time monitoring, and post-study data analysis.
The third principle is the theory-grounded customizability of interaction variables. Currently, the platform organizes supported interaction variables into four layers drawn from CSCW and HCI research: information flow, action structure, social framing, and system responsiveness. Each variable is bound to dedicated interface design options or controller specifications where researchers can independently manipulate it; for instance, they can manipulate the communication channels to enable private messages, broadcast messages, or disable messaging for all participants. The specific interface modules and controller settings that correspond to each layer are detailed in Section 3.2.
The fourth principle is native, standardized support for agent integration to ensure experimental validity and reproducibility. We designed a rigorous ACP to standardize agent integration for consistency and reproducibility. The protocol serves two high-level purposes: aligning agents’ perceptions and interactions with the environment to mimic the same experimental constraints as remote human participants, and maintaining consistency of these constraints across different agents, regardless of agent architectures. The protocol standardizes agents’ access to the environment, action spaces, communication channels, and the structure of all data sent to and from the agents, which also supports comparative evaluation of different agent architectures natively.
@@ 3.2. Architecture Overview
The platform architecture follows a client–server model (Figure 1). Participant interfaces, ACP, and the researcher interface connect to an experiment controller over a real-time channel. In particular, ACP ensures that agents can periodically perceive the experiment states and interact with the experiment environment in a consistent way. The controller manages experiment state, validates and commits actions, and synchronizes all logs with the data storage. Here, we present design and implementation details of each component.

### v2

@@ 2.3. Remote Human Collaboration as an Analytical Lens of Exploring Human-Agent Collaboration
A first set of insights concerns the limitations of low-bandwidth communication. In particular, Media Richness Theory explains why low-bandwidth channels carry less contextual information and make negotiation harder (Daft and Lengel, 1986), with several experiments demonstrating that “leaner” media like text are less effective than “richer” media like video for conveying ambiguity and building consensus (Short et al., 1976; Straus, 1997). These results are critical to human-agent collaboration, where text is the dominant modality and interface cues such as typing indicators or meta-messages can meaningfully alter users’ interpretations of an agent’s intent (Erickson and Kellogg, 2000; Iftikhar et al., 2023).
@@
A second challenge involves the “mutual knowledge problem,” where collaborators lack common ground and must infer what others know (Clark and Brennan, 1991; Cramton, 2001). Interfaces that add timely acknowledgments and visible states of collaborators could reduce these failures (Kraut et al., 2003). Research on workspace awareness further strengthened this idea by defining cues about presence, activity, and intent that help people predict collaborators’ actions (Dourish and Bellotti, 1992; Gutwin and Greenberg, 2002). For LLM agents, this can be achieved by asking agents to generate future plans with rationales visible to human collaborators through interface designs.
@@
Task structure and coupling also shape distributed collaboration. Distance complicates timing, handoffs, and coordination under partial information (Olson and Olson, 2000; Hinds and Bailey, 2003). In particular, a variety of lab-based experiments were designed to systematically analyze the impact of different interaction variables in a controlled setting. For instance, DayTrader manipulates the communication medium during negotiation in a social-dilemma setting and identified that the choice of medium shifts trust formation and deal quality (Bos et al., 2002). Bos et al. (2004) further illustrated trading biases and subgroup formation caused by the differences between in-group (collocated partners who can talk face-to-face) and out-group (remote partners) in the Shape Factory experiment. NeoCITIES imposes time pressure and role interdependence in a shared information space (McNeese et al., 2005; Schelble et al., 2022), whereas Passcode examines how teams form and update shared mental models during problem solving when only partial clues are available (Gero et al., 2020).
Although LLM agents differ from human partners in their ability to perceive the context or disclose uncertainty (Liao et al., 2020; Hoffman et al., 2023), these capabilities enable adapting classic human-human collaboration variables for human–agent settings. Our platform supports fine-grained manipulation of communication modalities, visibility of agent state, and task coupling within re-implemented experimental paradigms tailored to human–agent collaboration.
@@ 2.4. Research-Focused Toolkits and Platforms for Collaboration Studies
A wide range of research platforms support collaboration studies (Fig. 2), including classic groupware toolkits, online behavioral experiment platforms, game-based environments, and newer agent frameworks. While each family offers relevant capabilities, none provides the experimental control or human–agent parity needed for systematic studies of human–LLM collaboration.
Groupware and General-Purpose Behavioral Platforms. Early HCI toolkits such as Share and GroupKit (Greenberg, 1990; Roseman and Greenberg, 1996) enabled synchronous distributed tasks through shared views and real-time session management. More recent platforms (e.g., LabintheWild (Reinecke and Gajos, 2015), Empirica (Almaatouq et al., 2021), oTree (Chen et al., 2016), and nodeGame (Balietti, 2017)) provide scalable participant management, messaging, and data capture for human–human experiments. Nevertheless, these platforms focus only on human studies, offer no support for agent integration, and do not allow manipulation of collaboration variables through the interface.
Game-Based Research Environments. Game environments such as Overcooked (Carroll et al., 2019) and PairIt (Ju and Aral, 2025) support precise coordination and performance measurement, enabling controlled studies of teamwork dynamics. Yet their domain-specific mechanics limit structural adaptability: researchers cannot repurpose them to vary communication modalities, alter information coupling, or instantiate classic CSCW experiment paradigms.
@@
Agent-Oriented Frameworks and Social Simulators. Frameworks like AutoGen, AgentVerse, and CAMEL (Wu et al., 2024; Chen et al., 2023; Li et al., 2023) support multi-agent task automation and agent–agent communication. Others (e.g., WebArena (Zhou et al., 2023), Melting Pot (Leibo et al., 2021), and generative social simulacra (Park et al., 2023, 2024; Piao et al., 2025)) focus on agent performance, environmental fidelity, or large-scale population simulation. Nevertheless, these agent-oriented frameworks and simulation platforms emphasize agent performance, simulation fidelity, or agent–agent dynamics, not for human-agent interactions.
2.4.1. Comparison with Existing Research Platforms
Table 1 presents a systematic comparison of existing behavioral experiment platforms and agent frameworks that are relevant to ours in terms of dimensions that are critical for human-agent collaboration.
Platforms like oTree (Chen et al., 2016) and Empirica (Almaatouq et al., 2021) have set the standard for reproducible human-subject research by providing robust session management, experiment logic, and data logging. However, these frameworks were oriented for Human-Human interaction. Integrating LLM agents into these platforms requires significant ad-hoc engineering (e.g., setting up separate WebSocket servers to handle agent streaming), and there is no platform-level protocol that ensures agents perceive and act under the same constraints as human participants. In our system, the Agent Context Protocol (ACP) constructs an experiment-specific context from the configuration, issues periodic state summaries governed by a perception interval, validates agent actions against schemas and policies with error handling, and commits accepted actions atomically. These semantics create parity between agents and remote humans.
Conversely, frameworks like AutoGen (Wu et al., 2024) are designed to enhance agents’ automated task-solving capabilities, which is contrary to the goal of prioritizing experimental manipulation of interaction variables for human–agent collaboration. In a research study, a scientist may need to intentionally degrade communication (e.g., ”disable chat”), introduce latency, or hide information to test a hypothesis about reliance. In contrast, our platform prioritizes Interaction Control over execution speed or standalone task-solving accuracy. Through the Interaction Control Interface (Section 3.3), researchers can systematically manipulate the information flow and action space available to the agent, enabling valid comparative studies that isolate specific collaboration variables.
Finally, environments like Overcooked-AI (Carroll et al., 2019), CREW (Zhang et al., 2024a), and Pairit (Ju and Aral, 2025) provide rich testbeds for a fixed domain-specific task but have limited structural adaptability for research questions that require different coordination structures or staged procedures. Instead, ECL and the controller in our platform allow researchers to reconfigure objects, action schemas, and views to implement both classic and new experiments. We illustrate this by conducting evaluation experiments in Shape Factory and Hidden Profile, and adapting DayTrader, Essay Ranking, and Passcode with only changes to configuration and view binding, as shown in Section 3.4.
In summary, our system combines (i) a declarative configuration model that separates state, policy, and views to support controlled interface manipulations, and (ii) a protocol that gives LLM agents the same perception and action semantics as human participants, with controller-side validation and atomic commits. These design choices enable controlled comparisons in human–agent collaboration that are difficult to realize in general multi-agent frameworks or monolithic behavioral platforms.
Table 1. Comparison with behavioral and agent frameworks on dimensions relevant to human–agent collaboration research.
Table 1. This table shows a systematic comparison with behavioral and agent frameworks on dimensions relevant to controlled human–agent collaboration research. This table compares the authors’ platform with AutoGen, oTree/Empirica, overcookedAI/CREW in different feature dimensions.
Feature Dimension Our Platform AutoGen (Wu et al., 2024) oTree / Empirica (Chen et al., 2016; Almaatouq et al., 2021) Overcooked-AI / CREW (Carroll et al., 2019; Zhang et al., 2024a)
Primary Focus
Controlled human–agent
Collaboration experiments
Multi-agent task pipelines
Behavioral experiments
with humans
Coordination in a fixed game
Configuration Model
Declarative ECL
(objects/actions/policies/views)
Programmatic
orchestration
Programmatic study
scripts
Fixed environment / Gym-style API
Agent Integration
Standardized protocol
(ACP) for parity
Native or ad-hoc message
passing
No native agent
semantics
Built-in agents or RL policies
Experimental Control
UI/media toggles, timing,
blinding, variable exposure
Low; optimized for
completion pipelines
Randomization/payoffs;
no agent-facing controls
Fixed rules; limited factor control
Architecture Pattern
Decoupled controller/view;
state–view separation
Tightly coupled
orchestration and execution
Monolithic web
framework
Game engine simulation
Example Experiment Fit
Shape Factory, Passcode,
Hidden Profile variants
Code generation, tool use,
web tasks
Public-goods, auctions
dictator games
Overcooked variants
@@ 3. System Design
To address the need for reproducible and extensible research on human-agent collaboration, we designed and implemented an open-source, fully configurable research platform. Unlike existing platforms that either focus on human-human experiments (e.g., Empirica (Almaatouq et al., 2021), oTree (Chen et al., 2016)) or development of autonomous agents (e.g., AutoGen (Wu et al., 2024)), our platform serves as a flexible infrastructure that brings LLM agents into the controlled experimental environments traditionally reserved for human collaboration studies.
The platform’s architecture encompasses four key components: Participant Interface renders the interaction environments for participants, Researcher Interface allows for the configuration of experimental variables, Experiment Controller that serves as the backend engine, and an Agent Context Protocol (ACP), a novel integration layer that standardizes how agents perceive and act.
Researchers can adapt our platform for diverse experimental paradigms in two ways: through a declarative Experiment Configuration Language (ECL) (example provided in Appendix B) that defines all entities, actions, and interface modules, or by directly extending the open-source codebase. By defining entities, actions, and interface modules through ECL rather than ad-hoc code, our platform supports seamless adaptation of existing paradigms, such as social dilemma games (Bos et al., 2002) and collaborative decision-making tasks (Zheng et al., 2023a), and the development of entirely new ones. Researchers can adjust experiment parameters and manipulate interaction controls for different conditions via the researcher interface.
We address open-science needs by releasing source code and by encouraging the sharing of configuration templates, expecting our platform to support the broader research community in human-agent collaboration research.
@@ 3.1. Design Principles and Rationale
The platform targets synchronous, interface-mediated collaboration among mixed human-agent teams. We articulate four design principles (DP) grounded in HCI and CSCW theory: enabling the replicability of classic paradigms, translating collaboration theory into testable controls, ensuring agent-human parity, and synchronized process logging for post-study analysis.
3.1.1. DP1. Generality via Abstraction.
Existing research platforms, as shown in Section 2.4.1, are often designed or hard-coded for specific tasks (e.g., text-based conversation only), which largely limits their generality. To support a broad range of collaboration research, our research platform must be easily adaptable across different experimental paradigms without requiring redesign or re-implementation of the core architecture. To address this challenge, our first principle focuses on achieving generality through both logic and interface abstractions to ensure its paradigm adaptability.
For logic abstraction, the platform must abstract the fundamental interaction mechanics that recur across paradigms (e.g., resource distribution, turn-taking, and information asymmetry) away from the specific experimental content. The Experiment Configuration Language (ECL) allows researchers to either reuse existing classes or instantiate new classes, define custom attributes, and bind them to experiment-specific artifacts and visual elements, correspondingly. For interface abstraction, we deconstruct the participant interface into visual modules (e.g., the Social Module) that can be customized independently. As a result, researchers can seamlessly replicate classic experiment paradigms or rapidly prototype entirely new human-agent collaboration tasks by altering the interface design, interaction flows, and available information for each participant to fit specific experimental needs.
3.1.2. DP2: Translate Collaboration Theories into Testable Controls
Decades of HCI/CSCW research have produced profound knowledge on human collaboration; nevertheless, these high-level findings have not been tested in the emerging context of human-agent collaboration. For instance, prior research shows that variables such as workspace awareness, media richness, and social signaling fundamentally shape team dynamics. In principle, these high-level collaboration theories have to be translated into concrete, manipulable variables or feature designs for empirical testing. Thus, our second principle is to turn these theoretical findings for collaboration into modular and configurable Interaction Controls.
Specifically, the platform’s research interface includes controls grounded in established collaboration theories, enabling researchers to toggle, restrict, or customize the corresponding visual modules (e.g., private inventory, shared dashboards). Researchers can thus isolate independent variables (e.g., customizing the ”Social Dashboard” to test grounding constraints) and measure their specific effects on human–agent collaboration. This principle ensures that the platform’s configurations are meticulously designed to serve as valid independent variables for hypothesis-driven research.
3.1.3. DP3: Reproducibility via Standardized Human-Agent Parity
A critical consideration in emerging human-agent collaboration is how to ensure the agents are integrated consistently. Specifically, the system must enforce that agents operate under the same perceptual and temporal constraints as remote human collaborators to ensure valid comparisons between human-human and human-agent teams. Besides, agent integration must follow a strict schema regardless of the underlying model architecture to ensure the validity and reproducibility of experiment results. Our third design principle directly addresses this concern by enforcing standardized agent integration into the platform by leveraging the Agent Context Protocol (ACP). The ACP filters system states into a human-analogous context window, standardizes agents’ access to the environment, action spaces, and communication channels, and ensures that experimental results are attributable to the agent’s social and cognitive behavior rather than unequal system access.
3.1.4. DP4: Synchronized Process Logging
Investigating human-agent collaboration dynamics requires analyzing beyond the collaboration outcomes; instead, we should follow established research on remote human collaboration to emphasize the process of interactions. In human-agent collaboration, the process data is multimodal, encompassing observable human and agent behaviors as well as latent agent reasoning and planning trajectories. The fourth principle tackles this requirement by enforcing synchronized process logging. Specifically, the experiment controller aligns interface events, human messages and actions, and agent prompts and tool calls on a single, synchronized timeline. The process logging synchronization allows researchers to tie outcome differences to specific coordination moves, breakdowns, or repairs, which can support granular mixed-methods analysis of when an agent’s reasoning is misaligned with the situation or their human collaborators’ actions.
Figure 3. An example of the participant interface for the Shape Factory experiment (Bos et al., 2004). The interface consists of five fully configurable modules: My Status displays the participant’s private status, My Actions displays permitted actions, My Tasks displays the tasks to be completed, Social provides available interaction channels, and Information Dashboard shows other participants’ real-time information. The experiment’s timer is shown at the top.
Fig. 3. This figure shows the participant interface of the Shape Factory experiment. The interface is divided into five modules. ”My Status” shows each participant’s private resources and condition. ”My Actions” lists the actions the participant can take. ”My Tasks” displays current goals or assignments. ”Social” provides communication options with other participants. ”Information Dashboard” shows other participants’ real-time status. A session timer is positioned at the top.
@@ 3.2. Architecture Overview
The platform architecture follows a client–server model (Figure 1). Participant interfaces, ACP, and the researcher interface connect to an experiment controller over a real-time channel. In particular, ACP ensures that agents can periodically perceive the experiment states and interact with the experiment environment consistently. The controller manages experiment state, validates and commits actions, and synchronizes all logs with the data storage. Here, we present design and implementation details of each component.